For the given functions ff and gg , evaluate the given composite function.

f(x)=5x2+2x,  g(x)=x−9

(f∘g)(2)
Answer #1

5₂2 & 2n f(x) = g(x) = x-9 (fog)(2) = ? fog) (a) 512109] + 29 (2) 5 [n- 97² 2 (a-g 5/ x2 +81 - 18x] + 2n-18 5x² 405 gon 2n-18